a metal has a specific heat capacity of .899 Jg-1k-1 How much energy is required to raise the temperature of 2.00 Kg of the metal by 80 K


1
Expert's answer
2022-05-17T11:30:02-0400

Solution:

q = m × C × ΔT

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J K−1 g−1)

ΔT = change in temperature (K)

 

m = 2.00 kg × (1000 g / 1 kg) = 2000 g

C = 0.899 J K−1 g−1

ΔT = 80 K


Thus,

q = (2000 g) × (0.899 J K−1 g−1) × (80 K) = 143840 J = 143.84 kJ

q = 143.84 kJ


Answer: 143.84 kJ of energy is required